OpenStudy (anonymous):

f(x)=x^4 -5x^3 +9x^2 Find the intervals on which f is increasing, the intervals on which f is decreasing

OpenStudy (anonymous):

typo in the question is it \[f(x)=x^4 -5x^3-9x^2\] or \[f(x)=x^4 -5x^3+9x^2\]??

OpenStudy (anonymous):

The second one

OpenStudy (anonymous):

then the derivative is \[f'(x)=4x^3-15x^2+18x\] or \[f'(x)=x(4x^2-15x+18)\] if you are lucky the second part factors

OpenStudy (anonymous):

The second part does not factor, so that's my problem

OpenStudy (anonymous):

oh actually that makes you even luckier

OpenStudy (anonymous):

the second part \[4x^2-5x+18\] has no real zeros, so it is always positive and you can ignore it

OpenStudy (anonymous):

@satellite73 so when it does not factor it means that there are no real zeros?

OpenStudy (anonymous):

therefore the derivative is negative if \(x\) is, i.e. negative on \((-\infty, 0)\) and positive otherwise that make your function decreasing on \((-\infty,0)\) and increasing on \((0,\infty)\)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if it does not factor over the integers, you can always find the zeros using the quadratic formula, but if you use it in this case you will see the zeros are complex, so no real zeros

OpenStudy (dan815):



OpenStudy (anonymous):

So the stationary point would be (0,0) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@satellite73 the critical point would also be 0

OpenStudy (anonymous):

you can see it pretty clearly here http://www.wolframalpha.com/input/?i=x^4+-5x^3+%2B9x^2

OpenStudy (anonymous):

as far as i know, stationary point is a synonym for critical point

OpenStudy (anonymous):

@satellite73 my teacher didn't let us use the graphing calculator, she only let us use tables. Thanks for all the help :)